Want to learn what's involved in digging up an Ice Age Elephant? You've come to the right place!

Go to the Daily Log during the reconnaissance trip in June, 2000
Back to The Mastodon Project main page


What follows is a day-by-day account of the Hyde Park Mastodon excavation, written by those that were doing the work. It is not meant to be written in eloquent language, but the excitement often shines through. Read on! And if you have any questions about the dig, you can email any of the scientists that entered the information into the journal.

Sunday, August 20th- Pete N drives from Ithaca to Hyde Park, meeting Warren A. at the bus station in the early afternoon. Jim Sherpa and Bill Perks, and John and Vicki Chiment drive to Hyde Park and meet with the Loziers at the pond. Paul Harnik and Rob Ross rent a U-Haul in Ithaca, fill it with supplies and drive to Pleasant Valley where they meet up with the rest of the crew at the Pleasant Valley Motel. Dinner and sleep. (Paul Harnik - [PGH] - email him)

Monday, August 21st-
Early AM wake up to get to the pond, unload supplies, pick up additional supplies and start pumping. 4 pumps going and pond drains within a few hours. During the draining, Paul Harnik, Jim Sherpa, Pete Nester begin to experiment with the coring device, taking test cores in the wetlands to the north of the pond. Augering goes fairly smoothly and different bits are experimented with. For some reason in the lowlands to the north, the peat layer seems absent in the cores. During this time, Bill and John return to the site, as well as Elizabeth Humbert and Curt Banta. Student and faculty members from Mount Holyoke (Diane and others(?)) are also on site at this time. A core is taken in the lowlands immediately to the north of the pond. A series of skids are laid down on the drained pond bottom to access the peat islands in the southern portion of the pond. A core is taken on the northern most peat island of the two. Coring goes very well, and the hole bottoms out a little over 2m in a fairly coarse gravel layer-some cobbles larger than golf ball size. Decision is made to strip the peat off the 2 southern islands. This is begun, and 5 buckets from each island are collected. Decide to collect 5 more while this is being done, and @6:45 PM, Elizabeth Humbert finds the sacrum of the mastodon underneath the northern most island at the basal portion of the peat layer. Palpating the mud suggests more bones present - work done for day. Dinner and celebration at the Hyde Park Brewery. (PGH - email him)

Finally, success! Seconds after the first bone was uncovered by Elizabeth Humbert. The team from the Paleontological Research Institution include, from left to right: Warren Allmon, Curt Banta, Paul Harnik, Pete Nester, Elizabeth Humbert, Jim Sherpa. Photo taken by Robert Ross.

 

Tuesday, August 22nd-
Exploratory work continues in and around the bone zone revealed a fragmented portion of tusk, ankle bones, pelvic bone, scapula, and limb bone. Trenches are made by the front end loader on the most western side of the pond to collect high resolution and bulk samples of strata. Trenches are approximately 1m deep and strata exposed consists of peat and marl with some gravel interbeds. Nice wood fragments and mollusks in situ. After two trenches, this work ceases as strata is very liquified and a lot of water is beginning to flow into the bone zone. Newspapers on site this day. (PGH - email him)

Back to top

Wednesday, August 23rd-
Story runs in Poughkeepsie Journal on front page, Headline on newspaper box says "More mammoth fossils found!" Discovery Channel on site during the day. Auger is taken at southern end of the pond indicating some marl layers aside pond. No trenches made to avoid rising water level. Morning spent struggling with pumps which are clogging rapidly with clay and silt (and fish). Dominick's son (from the Discovery Channel) finds a limb bone in fill in the bank on east side of the pond. Work continues and Wayne (front-end loader operator) and Curt Banta work along bank towards the north, digging up fill and checking for bones. Core taken on northern island in center of the pond. Skull found in fill on the eastern side of the pond - upside-down - teeth present - seems in good shape. Lower jaw not present (yet). Work done to clean it - begins to rain - work stopped for day - fence set up. Elizabeth, Warren, Curt, Sara Bong and John Chiment leave for Ithaca. (PGH - email him)

Thursday, August 24th-
Cast of characters: Pete, Paul H., Rob and Jim S. from PRI and Mike, Lauren and Melissa (high-schoolers from Tom Lake's "Summer Scholars" program.) Arrived at site ~8AM. Pond was filled up again, but with 3 pumps running, it was pumped down in a matter of 3 hours. Paul H. uncovered a spruce (?) tree at far SE corner of pond. It was measured at ~20' long with branches still attached. Bone at far western edge of the "bone-zone" determined to be the other tusk. Paul H. worked from the spruce to the west, and uncovered little to nothing. (Peter Nester [PLN] - email him) ***

Friday, August 25th-
Arrived at site at ~7:45AM. U-Haul truck rented and buckets loaded in. Cast of characters: Jim S., Pete N., Paul H., Rob Ross, and Mike, Lauren and Melissa from Tom Lake's "Summer Scholars" program. With wooden pallets, a curved structure was built around the "bone zone" to try and keep the water out. Worked hard to divert drainage from the "zone" if at all possible. After this curved pallet structure was built, a toe-bone was found "outside" the structure by Pete as he was moving slop to help with drainage. It was determined that one of the long bones just inside of the "structure" was, in fact, one of the tusks, and it extended, unbroken, for at least 3' w/~20cm diameter. Rob and Jim left for Ithaca at about 2PM. Kathy Lozier helped in the pit, and towards the end of the day, the main effort was spent trying to move the sludge away from the "zone", working in the south and west directions. (PLN - email him)

Saturday, August 26th-
Discovery Channel again here filming at ~7:30AM. Warren, Jennifer Tegan, Pete arrive at ~7:15AM. John Chiment with about 12 Cornell students arrive from Ithaca at about 11:30M. Limit of western tusk determined. Retaining wall appears to be helping to keep water flow away. Main effort spent at south end of island. 5 partially articulated vertebrae found, maybe more soon, extending in SE direction. Hay bales used instead of pallets so that bones underneath are not disturbed. Continued working on skull, moving dirt from around, but leaving plenty of matrix around the skull for later excavation. Pete leaves at 4PM. (PLN - email him)

Set up "bulletin board" outside fence with laminated newspaper articles, fact sheet, bone chart showing what we've found so far, and PRI banner. Visitors seemed to like it and read it. Jen Tegan and Warren inventoried and reorganized bagged bones that have so far been extracted and stored in garage. Dwight Warren in most of the day - he pulled out his cast skull and Discovery Channel shot it. Tent put up over skull and it is covered with blanket and towels to keep it moist. Lower jaw found ~4PM. Discovery Channel did "line up" of Cornell University students to show how dirty they got. Found lots of artic thoracic vertebrae, large ribs and, at the end of the day, two long bones. Quit at ~6:30PM. Cookout for students at Lozier's - students slept in garage and tents. (Warren Allmon [WDA] - email him)

Sunday, August 27th-
Jen and Warren at site about 7AM to start three pumps. Warren went to get gas for pumps. CU students up and working by about 9AM. A team of three students worked w/John Chiment to chain saw and tie up a slice from the large spruce log at the bottom of pond. Extended green flag grid to the east. Discovery Channel filming again. Turning on and off pumps. Team of students worked on skull - working around edges and underneath. Finally stopped till we take it out. Rest of students worked on bailing out sludge from "pit" in center of pond. Sludge pile on west side of pond near steps has "crept" southward displacing main walkway. New superstructure built of pilings and 4 x 4's at north end of pit out from wall to allow walking over the shrinking "islands" of mud w/out boards in contact with bones. This took much of morning. Students quit at about 1PM and left at 2PM w/ John Chiment and Bill Perks. Warren's car died at pizza shop and he had to go rescue Jen in Larry's car. Pete arrived from Boston at about 5PM and took Jen and CU student back to Ithaca. Warren cleaned up and quit at about 6PM. Discovery Channel did standup interviews w/Warren and Larry Lozier and shot Sheryl in pit then did large piece of just Warren in pit. Also interviewed John Chiment. Lots of ribs found and uncovered today, as well as second scapula and more toes. Femur and tibia recognized (formerly thought to be second scapula). In general, it's difficult to be in the main bone bed without stepping on a bone. (WDA - email him)

Back to top

Monday, August 28th-
----no record----

Tuesday, August 29th-
Jim and Sarah arrived at 8AM. Spent most of the day with Volunteer Kevin and Cathy trying to develop drainage for "pit" area. Major situation in the pump area causing water to flow back into the "pit". Due to lack of volunteers we couldn't work efficiently enough to scoop out water. Same journalist from Poughkeepsie visited site. Talked to Jim about recent findings. Topo-surveyor came by survey site and areas around Lozier home. Few bone fragments (possibly skull) were found outside quadrants on the other side of the wood-wall. Wonder if this might be the spot where the skull originally was. (Sarah Bong [SLB] - email her)

Wednesday, August 30th-
11AM - Poughkeepsie photographer arrived (Spencer Ainsley). Monisha Ray arrived at 1:45 PM, she worked in the bone pit for about 2 hours, mostly bucketing up marl from the perimeter of known bone hot spots. She finds a caudal vertebrae in N1W1 #110. As mentioned above, the Poughkeepsie Journal photographer came. He took quite a few shots of the skull. It's overcast this morning, so Jim took off the canopy for better light, he used both a trowel and Eliza Sherpa with the skull for scale, also did a close-up of the teeth and a couple of shots of the tusk. Spencer expressed interest in photographing the ite professionally if we won't do it. Says he'll try to bring us a full set of photos taken on Friday or Saturday. To supply Spencer with the data, Sarah and Jim measured the tusk while it was uncovered - 7'6" along the median line. Following up on suggestion by Larry, Sarah and Jim dug a path from the western sump (spring hole) to the north sump. This works great and frees up a pump. A significant pirce of wood is exposed in the lower marl near the thin, fine gravel layer ay yje northern end of the connect ditch, we leave it in situ. We build a platform into the south end of the pond to the SE sump, the outflow goes down a piece of ply wood to prevent further silt build-up. Carefully placing the inflow hose we are able to nearly completely drain the bone pit. Although we have not cross-referenced with the map, the mere removal of all that water appears to have exposed new bones. (Jim Sherpa [JMS] - email him)

Thursday, August 31st-
Sarah, Jim and Eliza Sherpa arrive 8AM, we start the pumps and go get gas. Tom Lake arrives with about 6 volunteers 9AM. ALl sign in and sign waiver, for have to wait for the pond to drain. We rerun the bone hole pump-out feed out of the pond towards creek. The pump successfully runs the water up the gradual incline we've laid out. Soon, even the bone pit is fairly water-free. The intake in the bone pit needs frequent clearing and Jim discovers that going barefoot allows him to move easily about the pit without disturbing or stepping on bones. Keith (student volunteer) proves to be exceptionally helpful. Various other tourists and volunteers show up throughout the day. Jim finds what appears to be an ankle bone at the far eastern edge of N1E3 (on top of the marl) showing again the trend of small, stray bones to the E and NE. About 3PM, Elizabeth arrives and Sarah leaves. I'm happy ELizabeth is here, sad Sarah is gone. 4:00, Larry comes home early, he helps reconfigure the platforms around the bone pit. Jim took photos of the bone pit today. (JMS - email him)

Friday, September 1st-
Jim, Eliza Sherpa and Elizabeth Humbert arrive at 8AM. We have trouble starting the northern pump, it takes an hour+ to get it going. Tom Lake arrives with a contingent of volunteering. One of Tom's volunteers manages to quiet the pit pump by wiring up the muffler. Barbara, an archeologist, starts to redraw the bone map 2x scale by quadrants. Kurt, a student of _____(?), stops by to talk but immediately starts working. He works till quitting time. Jim takes more photos of site from all sides. N.Y. Times reporter stops by and interviews Elizabeth, Jim and Larry. *** The NY Times reporter stays 2 hours. Wayne brings his backhow by for Saturday. NY Times photographer will stop by about noon on Saturday. (JMS - email him)

Back to top

Saturday, September 2nd-
Warren and Pete arrive at 7:00AM, Jim and Elizabeth at 8AM. Pump in center of pond had trouble starting- needed more oil. Wayne (backhoe) arrived at 9AM. Soundshields picked up at Rob's house (sound person from the Discovery Channel) to shield noise of pumps for cameras. 3 projects happening coincidentally; 1-describing stratigraphy from SW side of the "pit", 2-searching for more bones in every direction away from the "bone zone", and 3-constructing a box around theskull to transport it to the garage. In the aftrenoon, New York Times photographer present - story to run in tomorrow's "Metro" section. After several hours of "boxing" the skull (from 9AM to noon), it was determined that more work needed to be done on it. Mainly worked on by Jim Sherpa, Dwight Warren, and Warren Allmon. At 2PM, skull lifted and moved. Put onto palettes on top of trailer. There was much rejoicing. Dartmouth students arrived in the afternoon. Peter Rowell was presnt taking photographs. Wayne poked around to the north of the skull along the eastern edge of the pond, but nothing turned up. He deepened the sump before leaving. Narissa Russell and her student arrived late in the day to look at bones. Claudia Rowe, NY Times reporter, there. (PLN - email him)

Sunday, September 3rd-
Warren and Pete at site at 7AM to find it completely flooded - water from stream had backed up into north end of pond - water almost over banks all around. Warren and Pete started 2 pumps- third was under water in middle of pond. Then started building dam on north end. Dartmouth people arrived to help. Pete and Elizabeth both went swimming to pull floating wood, etc from middle of pond. Pump brought up from bottom with use of Larry's boat. Larry called Hyde Park fire department and they brought pump truck at about noon - 4:30 PM - they used 6" hose to pump about 1,000 gallons/minute and shoot it into trees on east side. Warren moved some hay bales from pond to stream to help stop siltation from runoff. Third pump repaired and got running again finally. Pumps ran until about 10PM. At about 7PM, we finally saw bones. Sheryl Sinkow arrived to take photos. Shirley Egan arrived but spent most of day sitting and watching with nothing to do. We all left for motel at about 7:30 then all went to Valerie Furey's house for dinner. Jim Sherpa built out some sidewalks. John Chiment and Bill Perks arrived about 1PM with group from Cornell - not much for them to do - lots of sitting around and waiting for the water to go down. Pete cleaned out the garage and reorganized. Pete and Elizabeth left at about 4:30PM. Nerissa Russell and 2 kids left at about 1PM. Valerie and Pam Loughmiller did some digging on east side but found nothing - then they cleaned up coring equipment. (WDA - email him)

Monday, September 4th-
We were supposed to be in the New York Times today, but weren't. Warren and Pam on site about 7:30. 3 pumps started. Then 1 died temporarily - Warren rented another pump at Home Depot - moved a pump down into pit and this proved to work well to drain main bone area. Bones exposed better than ever. Shirley Egan there again, working in pit until about 2PM. She went out and got gas twice for us. Sheryl Sinkow took lots of photos. Bonnie ______(?), archeologist, stopped by. Jill Schneiderman brought about 40 students and about 4 faculty from Vassar at about 1PM. They replaced volunteers in pit and made lots of progress. Jim Sherpa left at about 1PM. Paul Harnik and David Kim arrived at 4:30. Warren, Pam L., Valerie F. left for Ithaca at about 5:30PM. (WDA - email him)

Tuesday, September 5th, 2000-
Pumped beginning at 0800. At noon, we were able to begin work. Water level might have been affected by shutting pumps down around 5PM of previous day. Jim Kennedy found three bones working north of the tusk (2 foot bones and a tusk fragment). Russ Fuller found a toe bone east of the pelvis. Paul H. spent some time with the news crew from NBC. Hyde Park weekly paper reporter came by and Paul spoke with them. NPR interviewed the Loziers by phone. We worked the later part of the afternoon, improving drainage. Diverted the spring to run towards the pump sump. Day was clear and temperatures were in the 60's. Wrapped up at 6:15PM. (Russ Fuller)

Back to top

Wednesday, September 6th, 2000-
Pumping began at 4:15AM. At 6AM, NBC crew arrived as well as John Chiment and Bill Perks with a truckload of buckets. Pond was drained (mostly) by 7:30 AM. David Kim and Paul H. cleaned the bones off between 7:30 - 8AM. The Today Show went off without any problems with all bones exposed and Larry and Sheryl doing a fine job. After TV crew left and Chiment and Perks as well bones were re-covered at 9AM and then covered with a tarp. Russ Fuller and Kevin (Poughkeepsie) volunteered for a day. We continued to expand the bone pit and did not turn up any new bones. Some new springs were uncovered, though the pump seems to be able to keep up. Russ spent a bit of time around the fragmented tusk and felt a few medium-sized (larger than your fist) fragments. Decided to leave them in place and put a tarp over them to stabilize them. At about 4PM, David and Paul almost collapsed from exhaustion - cleaned up the site and then sat on the banks and ran the pumps for another couple of hours. (PGH - email him)

Thursday, September 7th, 2000-
Started pumps at 7:30 AM. Just Paul H. and David until around 12:30PM. Moving slower today after the long day yesterday. Adrian Epstein volunteered all of the afternoon. Paul Harnik found some interesting pieces of wood, including a small branch with beaver marks on it. Pete Nester arrived around 5PM. Paul H. and David Kim return to Ithaca. (David Kim)

Friday, September 8th, 2000-
Pete arrives at 7:45, begins 3 pumps. 4th pump started at 8:30. By 10:15, bone zone clear enough to start working. Adrian filled buckets with peat from tarps - no lids on buckets. Continued to work the south wall - no bones found during the day. Another attempt at detailed stratigraphy was done along south wall. (PLN - email him)

Saturday, September 9th-
Three pumps running at 7:15. Fourth pump at 8AM - spark plug stripped, changed out for a different pump later in the day. Crowd from Mount Holyoke and Vassar worked in the pit, making more progress towards south and west wall. Island centered arouns SE corner of S1W1 was eventually entirely removed - matrix saved and bucketed (still no lids), trying to keep the peat from the marl and the marl from the peat. Rains came just before noon - lasted steadily until 3PM. At 3PM, began preparations to remove fractured tusk. It was determined that distal 4' of tusk was unbroken, but went underneath northern retaining wall. Extention built around tusk, and part of wall cut out to allow access to the distal portion of the tusk. Tusk was "pedastalled", and the exposed top 2/3 of circumference wrapped in paper towel, then newspaper, then it was "cable-tied", then wrapped in aluminum foil, then 2 layers of fiberglass cast applied to entire 2/3 circumference. Splints were applied to the top and sides of tusk before being wrapped in fiberglass. 18" lengths of 1"x4"'s were placed and driven beneath tusk in 3 locations, then cable ties wrapped around tusk, then rope wrapped around tusk in 3 locations. 3 people at tusk level and 3 people above, holding rope, pulled tusk out of ground. Weight estimated at 400-500 pounds (***this was later determined to be an over estimate. It only seemed heavy due to the awkward nature of the tusk and the slippery nature of the platforms, etc...[PLN]***). 5 people carried tusk up to walkway (slippery), and as it was being carried up, cracked about 30" from the tip. Casing failed - determined that at least 5 coats of fiberglass was needed. (PLN - email him)

Sunday, September 10th-
Pumps drained pond at 10AM. Pete and Warren on site. Narissa Russel, with assistance of Suzanne Gifford and Pam Loughmiller, began mapping bones in pit, finishing pelvis. Last day of working towards the south and west of pit, still finding nothng. People Magazine photographers on site around noon - took pictures of Larry, Sheryl and daughter Laura Lozier in the pit. Tusk, which was extracted yesterday, covered in paper towels and a tarp in the garage. Mapping of bones reconveined at around 3PM until about 4:30. Well over 200 visitors today. John Chiment took top part of spruce log with him back to Ithaca to give to Carol Griggs. (PLN - email him)

Back to top

Monday, September 11th-
Pam Lougmiller, Warren and Pete arrive at site at 7:30. 3 pumps running - pond down by 10AM. 4th pumps started to drain pit. Pete organized and catalogued bones in garage in the AM. Warren and Pam spent day taking horizontal locations of bones on eastern half of th epit, including pelvis and mandible. Two eastern quadrants completed. Fourth pump (north end) never ran correctly all day - Grand Rental coming out in the morning of 9/12 to service pump. Only three people there to work all day. Warren picked up two shelves for placing bones in garage - Pam assembled one. Tagging of bones in pit began, starting at #300. About 1/2 of them tagged. Janee Ansley and Elizabeth Humbert arrived at 4:30. Elizabeth and Pete stayed, Janee, Warren and Pam back to Ithaca. Elizabeth found Pete's missing shoe at 6:30 in the sludge along the north end of the pond. Generally a quiet day of progress. (PLN - email him)

Tuesday, September 12th-
Pete and Elizabeth arrive at the site at 8AM. 3 pumps started no problem, including pump at north end of the pond that was giving us problems the day before. Pond seemed to drain relatively well - down yb 10:30. 2 Vassar students at 10:30 and 12, helping to map bones in quad N1E1. Stake 0-0 appears to be off relative to rest of square - we set up another 0-0 stake about 25 cm to the north - this is now the 0-0 stake in terms of the bone map. Yesterday's map (as well as the day before) was keyed off of the line between stakes 0N1E and 1N1E. Today's map keyed off of 0N0E and 0N1E. Most of quad 1N1E completed today - mapping started by Pete with Elizabeth and 2 Vassar students in the morning, and continued in earnest in the afternon by Elizabeth (mapping) and 2 Vassar students (Josh and Holly) in the pit, taking measurements. Sue Johnson from Drexel University arrived around noon - PhD student studying sexual dimorphism in mastodons. She took measurements of the pelvis, determining it to be a female. She also helped with the field measurements of bones. Skippy (Vassar student) built a second shelf in the garage for bone storage. Many teachers from Hyde Park there in the afternoon, gathering materials (mostly video) to show students. First thing in the morning (whole pond drained), bones organized and placed on shelf that was built yesterday in the garage. Wrapped up at about 5PM. (PLN - email him)

Wednesday, September 13th, 2000-
More mapping today. Pete and Elizabeth arrive on site at 7:15, pond drained by 10AM. During draining, new jacob staff (2m) formed for mapping. Sue Johnson here at 10AM, horizontal plane mapping by Elizabeth in quad N1E1 w.Sue and Molly (from Vassar). Quad predominantly finished (except for a few small bones). Began quad S1E1 in afternoon. Peter with aid of Vassar students taking vertical measurements from datum pole. Dutcbess CC students (Celia and Dan) helped move pallets back from retaining wall. It was being undermined and collapsing. Tusk in precarious position. Mudbank it is on is being work away - need to remove soon. Last bone tagged is #373. Sara arrived at 5:30. Elizabeth G. arrived at 4PM. Warren and Dan Fisher arrive tonight. 3 rooms reserved at Econo Lodge. Pete and Elizabeth return to Ithaca. (PLN - email him)

Thursday, September 14th, 2000-
WDA, Sarah Bong and Dan Fisher on site at 7AM. Started pumps with no problem. ELizabeth G. there at about 10. High school class from Darien, CT arrived at about 10AM. Wayne Mallen brought Boy Scout troop by in the afternoon - several parents too. Some scouts going from their geology merit badge. Warren gave tours and ran errands much of day. Dan, Sarah and Elizabeth G. worked on mapping, plastic tagging of bones rest of day. Dan Fisher set up his bipod for aerial photos of bone pit and worked with Vassar students most of afternoon - took a roll of slides for stereo photos. Jill Schneiderman stopped by. Bonnie Bogamil worked with map crew in late afternoon. Susan Johnson and Jay Tinker volunteered all day. Jim Sherpa arrived at around 2PM. Dan Fisher and team in pit tagging bones until 7PM. Rebecca Rothbaum from Poughkeepsie Journal and Kathleen Kelley from People Magazine at site. (WDA - email him)

Back to top

Friday, September 15th, 2000-
Article on front page of Poughkeepsie Journal today. Rain last night but water not seriously higher. (WDA - email him)

Saturday, September 16th-
Plan to excavate larger bones today. Jim, Dan Fisher and Rob arrive at 6:30 AM to pump pond. Dwight Warren shows up soon after. Discovery Channel people arrive about 7:45. By 9:15, crowd has developed. Photographer, Jay Tinker, others on site. At 9:30, there are 40+ on sides, 18 people "touring" in pit, and the crane just showed up. Most people are just observers.

Post-day wrap up- An incredible day! A remarkable mixture of pandemonium and chaos, of professionalism and standing around. Brief: Around 9:30-1:30, used "cherry-picker" with bucket to take overhead shots. First our photographer-a volunteer-taking 35mm shots, then Discovery (Dominic) taking film and still digital images of the pit empty, then Dominic filming Dan in the pit, then Steve Westin and Pam Loughmiller taking digital images for possible stereoscopic images. Huge number of volunteers and random visitors, watching from driveway and edge of pond, which was roped off. At one point, Tom Lake counted 78 cars up and down the road; total people unknown-a couple hundred during day(?). Lunch-12 pizzas from Mona Lisa's-came at 2PM. A soda machine came from Larry's brother (in law?). At about 2:30, had a volunteer meeting and organized into teams with team leaders-Dan in pit, Rob in documentation, Dwight Warren in transporting bones up, John Chiment in tent curating bones as they arrived, Jim general manager, facilitating problems as they arose. Each of us had a group of people to give jobs to. Selecting volunteers for teams chaotic, but it seemed to work itself out. Dave Burney was in the pit with Dan-big help in sedimentological interpretation, and also in use of pumps (with which he has extensive experience through his own research.) Jennifer _____(?) took digital photos as bones were taken out, Suzanne Gifford marked bones onto digital photos she'd taken and had printed just before work begun, Rob took notes dictated by Dan. Discovery Channel continued to film short segments and ask questions to Dan, sometimes doing 3 or 4 takes on the same answer. Crowds remained to watch-cheer went up when first bone came out of the pit at about 3PM, and there was lots of excitement about removing the mandible. At 5:30, bone collection stopped; 13 left ribs, 8 right ribs, 8 vertebrae, 5 left fore foot bones, 1 right fore foot bone, a couple parts of the sternum, left ulna, left radius, left scapula (40 bones) had been removed. Dan accounted for bones, volunteers wrapped bones in bubble wrap and plastic bags, and stored them in the garage. (Robert Ross [RMR] - email him)

Sunday, September 17th-
Second day of bone removal. Arrived at site at 6:30AM-Jim, Rob, and Dan. We were on the front page of the Poughkeepsie Journal this morning-aerial view of the site. Started pumps-one pump shut off after a while because of water in the pump. Dwight Warren showed up and was able to drain tank and add oil, and got it started. Otherwsie a quiet start to the day. Al Guyro, Discovery Channel producer, came. John Chiment and Bill Perks showed up with about 20 Cornell students about 9:45, as well as Warren. Public started coming at 10AM. Excavated, with Dan leading, until 4:45. Got most of work done, with a huge hurry in last hour, especially after an articulated hind foot was found with more than 20 bones. Jim Sherpa and Dave Burney also in pit. Finished removing exposed bones about 5:30. Finished cleaning up by about 7:30. (RMR - email him)

Back to top

Monday, September 18th, 2000-
Spend morning finishing removing the lst few bones and documenting them. Some volunteers helped screen matrix from beneath the bones. Continued to remove marl at the base of the pit to make sure we aren't missing anything. Found one small bone fragment near pelvic area. Some volunteers worked on wrapping (and rewrapping) some of the bones. Jim began to take tusk measurements. PRI staff on board that day included Jim Sherpa, Paul Harnik, and David Kim. At least 8 volunteers assisted. (PGH - email him)

Tuesday, September 19th, 2000-
Jim continued tusk measurements and designs for its removal. Paul and David with a group of volunteers continued matrix removal and screening. About mid-day, it began to rain. Cold rain off and on for the rest of the afternoon and into the night. No new bones or fragments turned up. Beginning to hit a dark blue-gray clay in parts of the pit. (PGH - email him)

Wednesday, September 20th, 2000-
No rain today. Paul and Dave and Patrick ____(?) continue to mine the pit with the help of various volunteers. Jim works on box construction and talks to Dan Fisher about tusk removal. Slow but steady stream of people stop by the site to just look. Patrick proves once more to be a star volunteer. Paul H. and Dave Kim leave for home. (JMS - email him)

Thursday, September 21st, 2000-
Jim here alone, but only till the pond pumps down. Fair weather. A steady stream of visitors cause continual distractions. "We know there's nothing to see, but we want to look anyway," seems to be the sentiment of the day. Elizabeth Humbert arrives at around 2:30 PM and alleviates the outreach pressure. Jim speaks some more with Dan Fisher over the phone concerning the tusk. The pit goes on. (JMS - email him)

Friday, September 22nd, 2000-
Elizabeth and Jim onsite. Chilly morning-party cloudy. Began day at 7:30-all pumps working fine. Did not pump down bone yard in order to keep tusk wet. Mother (from Valkill-E. Roosevelt home) and son arrived at 10AM. They licked leaves out of buckets of matrix that had fallen in and, with the help of a group from DCC (environmental science class) began to put plastic wrap around tops of buckets. An eighth grade class that included Peter Rowell's son Zach arrived and toured the site. They aided in going through matrix and finding organics, etc. A group from Bard College arrived to tour the site. Kindergarten class from Montessori school arrived to tour the site as well. Also students from Marist (intro geology class). Volunteers dug out large spruce tree-tree is now pedestalled and 100% exposed. Jim and Elizabeth continued measurements for tusk box. Group from environmental science research (Paul Harnik's friends) helped dig out tree and go through matrix. End of day, branches were sampled (disconnected from tree). Calls throughout the day to PRI (Warren, etc.) Dan Fisher, Bill PerksÉ.continuing discussion on tusk removal. To fiberglass or not to fiberglass, that is the question!! Final decision is "yes" (I think). Bill Perks may come down Sunday to take tusk back to Ithaca. We need to mix the butte var NOW! Have not heard from turtle cop about log truck guy. Christie Sobel (an artist from Ithaca) and family arrived to help. Susan Rottman arrived for the day. (Elizabeth Humbert [SEH] - email her)

Back to top

Saturday, September 23rd, 2000-
Jim and Elizabeth H. on site by 7:30-pond water high. Volunteer groups from SUNY Oneonta, Boston Museum of Science, and Mount Holyoke (with Diane). Morning spent draining pond-Elizabeth went to Home Depot to buy some lumber and other supplies for the tusk box and tree box. Volunteers continue to sort through matrix (marl). A few hairs (?) and organic bits found. Continuing measurements for the tusk box-bottom and sides being constructed by Jim and the volunteers. Continued excavation of log-still pedestalled but being undercut. Began checking area north and east of retaining wall (dam) for other bones. Not much found. End of day, we constructed platforms to south of tusk for fiberglass people-doctor from Elmira-same doctor who did the fiberglass for the Chemung Mastodon tusks. Pump issues! Pumps keep breaking vacuum seal due to low water levels. Picked up Pete Nester at Poughkeepsie train station at 9PM. (SEH - email her)

Sunday, September 24th-
Elizabeth and Pete at site at 6AM. Pull cord on 2nd east pump breaks. Pete and Elizabeth move pit pump to east sump. Pumps up and running-they go for gas. Jim S. arrives at about 7AM, starts to trouble-shoot tusk box construction and prepare for tusk removal. SUNY-Oneonta volunteers have spent night in garage, and are now put to work. 9AM Bill Perks arrives, he is just ahead of John Chiment, who brings with him Vicki (his wife), John (his son) and the osteologist from Elmira. Joe Fox, videographer for the Discovery Channel, arrives. The pond is almost ready, but not quite, so he gies for a bite to eat. Jim briefs on procedures and concerns and at 9:40, doctor ______ (?) commences applying cast, Pete, Elizabeth and John assist as all else watch. Joe Fox arrives back just in time to film. The casting goes perfectly, at 11AM it is done. Before the tusk is removed to the box (casket, as everyone starts to call it), Jim, Pete and Elizabeth recreate the removal of the femur, humerus, and ulna from the pit with the help of Mike, Dwight and Sheryl. Miraculously, Jim finds an articulated, fleshed out 6 inch mammoth! Wow! (Hope Al gets a kick out of that.) The doctor has hung around, he has brought his two sons, and they get a cameo with the ulna. The "casket" is brought down to the pit and with the help of three volunteers from Oneonta, Joh Chiment (the Younger), Mike,Jim and Pete organize and remove the tusk. Lots f bubble wrap later, the top is put on and the tusk, snug in its casket, is carried out without a hitch and loaded directly onto Bill's truck. Smooth!! Lunch-then a weigt test for the spruce log is organized. It is lifted by hand (or rather twelve hands-6 people cradling the tree with towels underneath) but it cracks and splits in several places. Volunteers leave for home, Jim, Pete and Elizabeth do damage control on the spruce. Pack up and leave. (JMS - email him)

Monday, September 25th, 2000-
I arrived at the site at 9:30, surprised to find Jim still here. Jim briefed me on this past week or so, and told me what needed to be done over the next couple of days. Jim takes off at around 10:30AM, so it's just me and the bones. It's a beautiful, clear fall day. Aside from the monotonous hum of the pumps, it's rather peaceful. I had a lot of time to think about how amazing this discovery is and all the amazingly dedicated people involved. A few people stopped by to check out what was going on or to get an update. People seemed fascinated by the "sea" of red and white muck-filled buckets! Hmmm??? I spent most of the day picking up around the site and in the garage, running the pumps and covering buckets. (SLB - email her)

Tuesday, September 26th, 2000-
Arrived at the site at around 7:45AM. Cold, wet and rainy today, not much action around here. Got the pumps running around 8AM. They are all functioning properly and despite the rain, the pond drained fairly quickly. I threw tarps over uncovered matrix and buckets and spent most of the day in the garage out of the rain. (SLB - email her)

Back to top

Wednesday, September 27th, 2000-
Pete, Elizabeth H., Paul H. and Rob arrive at the site at 6:30AM and start pumps. Larry stops by briefly. Wayne arrives at 9AM. Dave Burney, Guy Robinson (from Fordham University) arrive at nearly the same time. Trouble with pumps, and figuring out the proper balance of pumps and "bone zone" not clear of water until shortly after noon. Rob and Paul assisted Dave and Guy in the pit, making a detailed stratigraphic record of the SE face. PRI took 10cm samples and placed them in baggies-Fordham took 5cm samples for pollen analysis. Bulk samples also taken by PRI. Pete and Elizabeth worked on the skull in the morning, trying to pad it within its crate for transport. Rob and Paul and Pete and Elizabeth all helped with placing wood at various points along the crate so that inner-tubes, inflated in place, would have a surface to brace against. U-haul was picked up at noon. At around 2PM, Wayne lifed the skull into Bill Perks's truck-the caret just barely fit. Styrofoam (instead of inner tubes) was placed beneath the skull for cushioning support. Bill Perks leaves for Ithaca at 3PM. Remaining b ones in garage wrapped in bubble wrap. Wayne trenched NE side of pond at 4PM for more sampling, photographing and descrbing. Pete and Elizabeth worked on the spruce tree. 6 splints were applied with cable ties to each individual section (3 1/2', 6', 5') of tree. Broken zones were sawed to make removal cleaner and easier. Dave and Guy finished and left at around 7PM. At this point, Pete, Rob, Elizabeth and Paul lifted lightest (and shortest) log onto a palette, which Larry attached to a tractor with a rope. This was pulled out along the southern bank of the pond and covered with a tarp. The second (and heaviest) was put onto same palette, but palette broke as it was being pulled up by Larry. Darkness fell, and Larry and Sheryl turned on garage lights as well as bringing out a stand of lights and shining them into the pond. Eventually, the two sections of the tree in the pond were hauled out by hand up the southern bank in 2' "heaves". All in all, a 14-hour workday. (PLN - email him)

Thursday, September 28th, 2000-
Paul H., Elizabeth and Rob at site at 6AM. Wrapped sections of tree in tarp and the four of us lifted spruce into U-Haul. Rob and Paul left at 7AM for Ithaca. Pete and Elizabeth began organizing garage and cleaning pumps. 2 pumps loaded into U-Haul and returned-2 pumps remained. Final story ran this morning in Poughkeepsie Journal. Picked up Fed-Ex from Warren and had Sheryl sign papers at 10:30AM. Returned to site to load all the bones into back of U-Haul. Went to Larry's work at 1PM, had him sign papers. *** On the road at 2PM. Returned to Ithaca at 6PM to PRI, and unloaded bones with the help of Will Ober, Warren, Jim S., Paul Krohn and about 4 volunteers (including Tracy Rozelle). Earlier in the afternoon, Newscenter 7 was there to film the skull being delivered by Bill Perks (it was in his driveway in Horseheads, NY all night) to PRI. (PLN - email him)

Saturday, September 30th, 2000-
John Chiment on site with volunteers from Cornell to bucket some matrix. (PLN - email him)

Wednesday, October 4th, 2000-
Pete and Elizabeth on site at 9AM. A third pump delivered from Grand Rental at 11AM. With 3 pumps working, pond, which was full at the beginning of the day, was drained by 4PM. While lake is still full, we moved floating wood to side of lake. Retaining wall was even found floating in parts, which made out job considerably easier. Removed styrofoam soundshields from pond and loaded them onto U-Haul. Cleaned garage - loaded all samples into U-Haul 1st, then styrofoam sound shields, then equipment. Garage swept. Several pump hoses rinsed and wrapped. Garage entirely cleaned. Left site at 6PM for hotel, then home. (PLN - email him)

Tuesday, October 10th, 2000-
Pete, Elizabeth and Paul H. leave PRI at 8AM, get to Hyde Park at noon. Rent 2 U-Hauls, pick them up, and after lunch, get to Lozier's pond at 1:30. Load about 300 buckets into both U-Hauls-much organic material and rainwater has fallen into these buckets-about half do not have lids. The other half only have plastic sheets over them, held on with twine. Buckets all loaded by 2:30PM. Next, scrap wood on SE and east side of pond is moved to common junk pile on SW side of pond to make Wayne Mallen's work easier (Wayne is scheduled to come to the site tomorrow and Thursday to regrade Larry's yard). Wood moved by about 3:30 PM-hit the road at 4PM. Back in Ithaca with 2 14' U-Hauls and Paul's car at 8:30 PM. Buckets unloaded the next day. (PLN - email him)

Back to top
Back to The Mastodon Project main page

Go back to the Daily Log during the reconnaissance trip in June, 2000


The Paleontological Research Institution
1259 Trumansburg Road
Ithaca, NY 14850 phone: 607-273-6623 fax: 607-273-6620
Questions about the Website? Tell us!